OpenStudy (anonymous):

Find the product. (y^5)^3 · y^9 A. y^24 B. y^17 C. y^60 D. y^42

OpenStudy (amistre64):

\[(b^n)^m=b^{mn}\] \[\large b^{mn}*b^p=b^{mn+p}\]

OpenStudy (amistre64):

*patent pending :)

jimthompson5910 (jim_thompson5910):

What is (y^5)^3?

OpenStudy (anonymous):

I wish I knew!

jimthompson5910 (jim_thompson5910):

(y^5)^3 = y^(5*3) = ???

jimthompson5910 (jim_thompson5910):

When you have an exponential expression raised to another exponent, you multiply the exponents

jimthompson5910 (jim_thompson5910):

So what is y^(5*3)

jimthompson5910 (jim_thompson5910):

y^(5*3) = y^15 so (y^5)^3 · y^9 y^15 · y^9 Now add the exponents to multiply these expressions y^(15+9) = y^24 So (y^5)^3 · y^9 = y^24